Kabbed:

Tammy is at the dentist's office waiting on her appointment. She notices that the 6-inch-long minute hand is rotating around the clock and marking off time like degrees on a unit circle. Part 1: How many radians does the minute hand move from 1:20 to 1:55? (Hint: Find the number of degrees per minute first.) Part 2: How far does the tip of the minute hand travel during that time? You must show all of your work.

Kabbed:

@Hero I have the first answer but I need help on the second part

Hero:

Just so we're on the same page, what did you get for the first part?

Hero:

@Kabbed

Kabbed:

sorry I thought i had the answer for part 1 but im still figuring it out

Hero:

Okay, would you like me to help you out with it?

Hero:

@Kabbed

Hero:

So the first step is to determine how many degrees there are per minute. To do this, divide the degrees in a circle by the number of minutes in an hour since one hour represents one revolution of the minute hand. In other words compute 360/60 = 6. So one minute is 6 degrees. Next, figure out how many minutes has transpired from 1:20 to 1:55. You are correct that it is 35 minutes. The angle between 1:20 and 1:55 can be found by multiplying 6 times 35 to get 210 degrees. Next convert the degrees to radians by multiplying 210 by \(\pi/180\). Finally, use the formula \(s = r\theta\) to find the the distance, \(s\), the minute hand moved. Remember \(r\) is the length of the minute hand.

Kabbed:

s = 6θ

Hero:

Looks good so far. By the way, when you find \(\theta\) it is standard to find the fractional expression, not the decimal expression.

Hero:

You can convert to decimal at the very end of your computation.

Kabbed:

what

Hero:

Have you attempted to find \(\theta\) yet? If so, post your work here.

Kabbed:

how do I find it

Hero:

You have to multiply the number of degrees between A (the minute hand at 1:20 ) to B (the minute hand at 1:55) by the conversion factor \(\pi/180\) to get \(\theta\)

Kabbed:

s = 60

Kabbed:

what is the symbol called

Hero:

How did you get "60" for s? explain

Kabbed:

0 is the symbol but I can't write the symbol down

Kabbed:

were trying to convert 3.6 radians to degrees

Kabbed:

3.6 * 180/pi

Hero:

\(\theta = 210 \times \dfrac{\pi}{180} = \dfrac{7 \pi}{6}\) \(s = r\theta = 6 \times \dfrac{7\pi}{6} = 7\pi\) now convert \(7\pi\) to decimal to get the approximate answer in inches.

Kabbed:

how do you convert

Hero:

with a calculator. Do you have one?

Kabbed:

yes okay

Hero:

Usually you would use 3.14 for pi if necessary

Kabbed:

21.991

Hero:

Finally

Kabbed:

i thought 210 * 180/pi was going to convert the degrees to radians but then it became 7pi/6 somehow or is 6 the radians

Hero:

The 21.99 approximates to an exact number

Kabbed:

so the tip of the minute hand travels 21.99 inches

Hero:

\(\dfrac{7\pi}{6}\) is the radian value of \(210^{\circ}\)

Kabbed:

okay

Hero:

21.99 can be approximated to an exact whole number value

Kabbed:

how does s = rtheta = 6 * 7pi/6 happen

Kabbed:

oh wait nevermind

Kabbed:

1:20 to 1:55 in degrees is 210 right

Hero:

Yes, but you still have not acknowledged my very last comment.

Kabbed:

so the answer is the exact whole number value

Hero:

I would round to the whole number. Do you know what that is?

Kabbed:

So the answer is 21

Hero:

No, it rounds UP to the next whole number.

Hero:

You've never rounded up before?

Kabbed:

22

Hero:

Yes. Unless it says to approximate the answer to two decimal places, I would just put the whole number as the answer. "22 inches" to be specific. Always include units.